I have a test coming up on Thursday and need help please show steps and what your doing this is really hard for me.

Find all real solutions of the equation (x-6)^2-36=0



Answer by nerdybill(7384)

Notice, I can rewrite:
(x-6)^2-36=0
as
(x-6)^2 - 6^2=0 (difference of squares)
.
factoring the left:
(x-6 - 6)(x-6 + 6) = 0
(x-12)(x) = 0
x = {0, 12}